satellite73 (satellite73):

the following numbers are the only numbers between 1 and 100 that have a special property. what is the property? 4, 6, 12, 30, 33, 36, 40, 45, 50, 54, 56, 60, 70, 81, 88, 90, 100

myininaya (myininaya):

the property is that they are btw 1 and 100

OpenStudy (anonymous):

only

myininaya (myininaya):

lol

OpenStudy (anonymous):

hint? write them out

myininaya (myininaya):

write 4,6,12,30,....50,54,...,90,100?

OpenStudy (anonymous):

four, six, etc

myininaya (myininaya):

four,six,twelve,thirty, thirty-three, thirty-six,forty,forty-five,fifty,fifty-four,fifty-six,sixty,seventy,eighty-one,eighty-eight,ninety,hundred

OpenStudy (anonymous):

another hint one hundred because "hundred" does not work

myininaya (myininaya):

four six twelve thirty thirty-three thirty-six forty forty-five fifty fifty-four fifty-six sixty seventy eighty-one eighty-eight ninety one hundred

OpenStudy (anonymous):

ok now what do they have in common?

myininaya (myininaya):

i give up

myininaya (myininaya):

do you think zarkon would know?

myininaya (myininaya):

lets test him

OpenStudy (anonymous):

I think there's nothing common in them...

OpenStudy (phi):

I have a decent idea what it is. Shall I post?

myininaya (myininaya):

yes

OpenStudy (anonymous):

you have started out correctly by writing down the words.

OpenStudy (phi):

The number divided by the number of letters when written out is an integer number. For example, 4/length(four)= 4/4= 1 6/length(six)= 6/3= 2 etc.

OpenStudy (anonymous):

nice.